OpenStudy (1018):

Hey guys, can someone explain the difference between these: -3^4, (-3)^4, 3^-4. Thanks!

OpenStudy (anonymous):

\[-3^4=-\left(3^4\right)=-(81)=-81\] \[(-3)^4=\bigg((-3)(-3)(-3)(-3)\bigg)=\bigg((3)(3)(3)(3)\bigg)=3^4=81\] \[3^{-4}=\frac{1}{3^4}=\frac{1}{81}\]

OpenStudy (1018):

Thanks!